OpenStudy (anonymous):

David is rowing a boat upstream. The river is flowing at a speed of 2 miles per hour. David starts rowing at a speed of 6 miles per hour, but as he gets tired his speed decreases (at a rate of 1 mile per hour, every hour). Which equation represents the speed of the boat for x hours spent rowing? y = x + 4 y = x − 4 y = 4 − x y = 6 − x y = 6 − 2x

OpenStudy (anonymous):

y = 4 - x is your answer here

OpenStudy (anonymous):

thank you /.\

OpenStudy (anonymous):

Given equations A and B as and , respectively, which expression will eliminate the variable x? A + B A − B A + B A − B A + B

OpenStudy (anonymous):

can u help on that ?

OpenStudy (anonymous):

he starts at 6 mph and the river is against him at 2 mph so 4 and then its - 1 per hour so make that x so ya y=4-x

OpenStudy (anonymous):

are there equations there in the blank spaces or no?

OpenStudy (anonymous):

no /.\ it didnt copy right one second

OpenStudy (anonymous):

Given equations A and B as 2/5x+y=12 and 5/2y-x=6 , respectively, which expression will eliminate the variable x?

OpenStudy (anonymous):

@dan815 did this one so im going to go with what he put. 5/2 A - B is your answer